OpenStudy (anonymous):

what is the circumference of a circular disc with a diameter of 10 cm?

OpenStudy (jhannybean):

The formula for circumference of a circle is : \(C = 2\pi r\) in order to find r, we take the diameter, and divide by 2. \(\large r=\frac{D}{2}\) What wouldyou get as your diameter?

OpenStudy (jhannybean):

radius*

OpenStudy (anonymous):

5 right?

OpenStudy (jhannybean):

Yes 5 is correct.

OpenStudy (jhannybean):

Now you just need to plug your values into your calculator :) \[C= 2 \cdot \pi \cdot 5\]

OpenStudy (jhannybean):

Do not round your value of \(\pi\) when doing so .

OpenStudy (jhannybean):

What do you get? :)

OpenStudy (anonymous):

31.4 is that correct?

OpenStudy (jhannybean):

Yes it is :) Good job!

OpenStudy (anonymous):

Thanks a lot :D

OpenStudy (jhannybean):

No problem :)